# Low-Acceleration Gravitational Anomaly from Bayesian 3D Modeling of Wide Binary Orbits: Methodology and Results with Gaia DR3

### v3

@@ 3.2 Consolidated gravity in a Newtonian regime
Figure 20 shows the results on the consolidated PDF of $\Gamma$ with the wide binaries in the Newtonian regime defined at the end of Section 3.1. The results with the thermal and flat priors on eccentricity ($e$) for the standard and alternative samples are shown. The results with the thermal prior agree remarkably well with the Newtonian expectation $\Gamma=0$, while those with the flat prior show a moderate bias towards the negative side. The bias with the flat prior is not surprising given that the implicit median of $e$ is biased, while at least some current data are not precise enough to strongly constrain $e$ regardless of the prior.
@@ 3.3 Consolidated gravity in transition and MOND regimes
Compared with the Newtonian regime, wide binaries in the transition ($0.500.15<s/r_{\rm{M}}<$ 0.50) and MOND ($2.860.50<s/r_{\rm{M}}<$ 2.86) regimes have smaller internal relative velocities, and the Gaia measurement uncertainties of RVs can be comparable to (or even smaller than) them. Also, the numbers of wide binaries are smaller than in the Newtonian regime. Therefore, larger statistical uncertainties of $\Gamma$ are expected. Moreover, as evident in Figure 12, the Gaia reported $|v_{r}|$ values are systematically biased upward (compared with the other precise velocity components) due to the relatively large uncertainties of RVs. The bias in $|v_{r}|$ means a (though smaller) bias in $v(=|\mathbf{v}|)$, which will induce a systematic bias in the consolidated PDF of $\Gamma$. This systematic bias will be estimated below using mock Newtonian samples.
@@
Second, the deviation is larger in the MOND regime than the transition regime, indicating a gradual deviation from Newton in the high-acceleration regime ($\gtrsim 100a_{0}$) to modified gravity in the low-acceleration regime ($\lesssim a_{0}$). This trend and the estimated value $\Gamma_{\rm{corrected}}=0.085\pm 0.040$ or $\gamma_{g}=1.48_{-0.23}^{+0.30}$ (after correcting for the bias described below) in the MOND regime based on the nominal sample agree well with the generic prediction $\Gamma\approx 0.07$ (or $\gamma_{g}\approx 1.4$) of MOND Lagrangian gravity models (e.g. Bekenstein & Milgrom 1984; Milgrom 2010). The value $\Gamma_{\rm{corrected}}=0.121\pm 0.049$ or $\gamma_{g}=1.75_{-0.32}^{+0.50}$ based on the alternative sample is also consistent with $\Gamma\approx 0.07$. These results through the new Bayesian methodology match well several recent results (Chae, 2023a, 2024a, 2024c; Hernandez, 2023; Hernandez et al., 2024a) based on various statistical analyses utilizing $v_{p}$ only.
@@
Now I describe the systematic bias in $\Gamma$ induced by the RV uncertainties. To estimate it, I use mock Newtonian binaries produced with the procedure described in Section 2.4. The reason for this is twofold. First, to estimate the bias in $\Gamma$, the unbiased value must be accurately known. While the unbiased value is exact (that is, zero) in Newtonian gravity, it is not so in any modified gravity. Second, based on all previous studies of wide binary gravity tests, any effective gravity experienced by wide binaries is not expected to be drastically different from Newton but rather pseudo-Newtonian or something similar. Thus, at least as a first-order approximation, Newtonian gravity can be used for the purpose of estimating the bias due to large uncertainties of RVs. However, I will also consider mock binaries produced with the observed Gaia binaries themselves by giving scatters to the observed RVs. It will be shown that the biases estimated with mock Newtonian and mock Gaia samples are similar. The nominal estimates will be based on mock Newtonian samples.
@@
Simulations using Gaia RVs themselves are also carried out in a very similar way. Here the Gaia reported RVs are used to produce mock RVs by adding scatters using the reported nominal uncertainties. Thus, in this case the uncorrected gravity based on the Gaia data provides the input gravity as opposed to Newtonian gravity in the above case. Because the input gravity is not known precisely in this case, there is a minor difference in the details of the procedure. That is to say, because the Gaia-reported RVs are already scattered quantities due to measurement errors, adding additional scatters can, in few cases, produce too wildly scattered RVs that lead to extremely unlikely PDFs of $\Gamma$, e.g. more than $4\sigma$ away from the “input” gravity (note that there was no such concern in the case of Newtonian simulation because RVs exactly follow the underlying gravity, or $\Gamma=0$, before adding scatters). Those extreme individuals are excluded when $\Gamma_{\rm{med}}$ is derived by consolidation in a mock sample.
Figure 25 shows the distributions of $\Gamma_{\rm{med}}$ from mock Gaia observations. The biases or shifts of $\Gamma$ are overall similar to the case of mock Newtonian observations shown in Figure 24, but show some differences in detail. Unlike the Newtonian case, the transition regime is displayed in the left column because some (though small) shifts occur. In the MOND regime, the shifts and the widths of the distributions are overall similar to those in the Newtonian case. The exception is the shift in the middle row. This appears to be due to the relatively large input value of $\Gamma$ while the median of the distribution is similar to the bottom row. In the transition+MOND regime, the widths are significantly narrower than in the Newtonian case, but the widths from the mock Newtonian and mock Gaia cases bracket the widths of the consolidated distributions shown in the right column of Figure 22. In line with Figure 22 and Figure 24, Figure 25 indicates that Newtonian gravity is inconsistent with the Gaia data, if any, even more strongly than in the other figures.
@@
So far all results are based on the nominal/standard inputs. Here, reasonable corrections or variations to the inputs are considered to estimate known (but ignored) systematics. They include the H2022 prior on eccentricities, perspective effects in observed velocity components, and Gaia FLAME masses. Note that perspective effects are calculated (see Appendix B) assuming that two stars in a pair are at the same distance, and thus I refer to them as first-order estimates. When the tiny distance difference is not measurable, it appears most reliable to assume that it is zero (see Appendix B for further). Note also that FLAME masses are available for about two thirds of stars, and thus nominal masses are replaced only for those stars.
@@ 4.1 Meaning of the results and an exploration of possible systematic errors
In this work, isolated (both in projection and line-of-sight) wide binaries are carefully selected (Section 2.3) and used for Bayesian 3D modeling of orbit and gravity. This work is based on two key assumptions that (1) the instantaneous relative motion between the stars in any observed isolated binary is at a phase of an unknown elliptical orbit and (2) the instantaneous motion can be modeled with pseudo-Newtonian gravity with a generalized effective gravitational constant $G_{\rm{eff}}$.
When the center-of-mass of a binary system orbits in the Galaxy, the relative motion between the two stars will suffer from perturbations due to long-range chance encounters with other stars and the Galactic tidal field (see Jiang & Tremaine 2010). Such perturbations can accumulate over a long period of time since the birth of the binary system (e.g. several giga years), leading to an evolved statistical property of binary populations as studied by Jiang & Tremaine (2010) in Newtonian gravity. The maximum projected separation for the wide binaries in the nominal sample (Table 1) is $\approx 25$ kau. This corresponds to $\approx 0.074r_{J}$ where $r_{J}\approx 1.65$ pc is the Jacobi radius for the median total mass of $1.82M_{\odot}$ based on Equation (43) of Jiang & Tremaine (2010). According to Newtonian simulation results shown in Figure 7 of Jiang & Tremaine (2010), presently observed binaries with $s/r_{J}\lesssim 0.074$ would be expected to follow Keplerian behavior well. This indicates that the assumption of elliptical orbits is valid as long as Newtonian gravity is tested with the wide binaries of this work. Thus, the disagreement (Figure 22) with Newton in low-acceleration regimes (along with the agreement in a strong-acceleration regime) is an unacceptable result in Newtonian gravity. Note also that the local dark matter density ($\lesssim 0.01M_{\odot}\text{ pc}^{-3}$: see Read (2014)) inferred assuming standard gravity cannot make any contribution to the gravitational anomaly.
However, in MOND or other modified gravity it is unclear how good the two assumptions will be and how to interpret the Bayesian inferred values of $\Gamma$ ($>0$) (i.e.  $G_{\rm{eff}}>G_{\rm{N}}$). In MOND gravity (e.g. Bekenstein & Milgrom 1984; Milgrom 2010), the rather strong external field $\approx 1.9a_{0}$ of the Galaxy can warp the orbit and the orbit may not be closed when the external field is inclined with respect to the rotation axis (Brada & Milgrom, 2000). In such gravity models, numerical simulations are needed to produce mock wide binaries and to infer $\Gamma$ in the instantaneous motions of the binaries. The values of $\Gamma$ can then be compared with those for the Gaia binaries.
@@
Whatever the case with MOND gravity, if at least some of these systems are pure binaries without kinematic contaminants or abnormal RV data, standard gravity will be ruled out more severely. Perhaps, it is unlikely that all of them have kinematic contaminants or abnormal RV data for several reasons. (1) All of them already satisfy the presently highest data qualities and 13 of the 16 stars have ruwe $<1.2$. (2) Gaia DR3 astrometric solutions are based on data collected over a time span of 34 months, and thus any hidden close binaries with periods of the same order or shorter have already been flagged and precluded. (3) The availability of the Gaia-measured RVs with relatively good precision indicates that the stars are likely to satisfy relatively good single-star solutions for spectroscopic observations. (4) The posterior PDFs of all parameters are well-behaved and the fitted values of $e$ are particularly reasonable. Therefore, it may well be that at least some of these binaries individually rule out standard gravity.
@@ 4.2 Wide binary results in the context of recent small-scale gravity tests
Isolated wide binary stars with separation more than several kau in the solar neighborhood are unique laboratories to probe low-acceleration gravity for two reasons. First, dynamics is simple because only two point-like objects are involved (if suspicious impure systems are properly rejected in advance) and their motions can be measured accurately. Second, wide binaries can test the basic nature of gravity generic to a broad class of nonstandard gravity models, that is, whether gravity is boosted as a function of internal acceleration and how much. If the Sun had a companion star (or an observable bright component) several kau away, the fictitious solar binary system would provide an ideal definite laboratory to directly probe gravity at low acceleration.
@@
Without a directly observable object at a distance of more than several kau from the Sun, the observed comets in a relatively inner part of the solar system have been considered to test gravity models by calculating (or extrapolating) their dynamical evolution up to large distances (e.g. Królikowska & Dybczyński 2020). For example, Vokrouhlický et al. (2024) have inferred the distributions of $1/a$ (or, binding energy where $a$ is the semi-major axis of the orbit) for hundreds of long-period comets in Newtonian and AQUAL models. This is an example of testing specific theories through modeling based on modeling assumptions including the MOND transition/interpolating function. In contrast, as stated above, the inferred value of $\Gamma$ from wide binaries is a pure measurement, based on movements observed directly at large separations and without reference to any nonstandard theories of gravity. In other words, motions in the MOND regime are directly caught in action rather than extrapolated.
@@
For certain theories of nonstandard gravitational dynamics, even planetary motions can be used to test them. Because planets are subject to high gravitational accelerations that are orders-of-magnitude higher than $a_{0}$, there is no gravitational boost in the sense of boosted acceleration or orbital velocity. [^10] As shown in this work and previous work, wide binaries also confirm that there is no gravitational boost for internal gravitational acceleration higher than $\approx 10^{2}a_{0}$. However, specific modified gravity theories such as AQUAL and QUMOND predict that the gravitational potential of the Sun becomes nonspherical even in the inner solar system due to the the external field effect of the Galaxy. This nonsphericity can be quantified by a quadrupole moment (Milgrom, 2009) of the potential.
@@
The quadrupole moment can be viewed as a second-tier gravitational anomaly that depends on the details of a specific model, whereas the acceleration-dependent gravitational anomaly such as the radial acceleration relation in galaxies (McGaugh et al., 2016) and the $s/r_{\rm{M}}$-dependent value of $\Gamma$ in wide binaries from this work is the first-tier anomaly that is generic to the MOND (modified gravity) paradigm. Hees et al. (2014) use 9 years of Cassini range and Doppler measurements to infer a value of the quadruple moment that is consistent with zero. The Hees et al. (2014) value of the quadrupole moment is used by Desmond et al. (2024) to test some MOND models and MOND transition/interpolating functions. As Desmond et al. (2024) demonstrate, the quadrupole moment can strongly constrain the MOND external field effect and/or the MOND interpolating function if the measured value of the quadrupole moment is confirmed and the MOND interpolating function can be well constrained in a wide acceleration range around $a_{0}$ by more accurate galactic rotation curves in the future.
@@
At the heart of the debate between standard gravity (with dark matter) and MOND or modified gravity (without dark matter) in the universe is whether a direct gravitational boost occurs at low acceleration $\lesssim a_{0}$. The low-acceleration gravitational boost is the fundamental nature that is shared by all models belonging to the nonstandard paradigm, while solar system tests discussed above are more about specifics of models. This means that even if solar system tests truly rule out certain modified gravity models, they cannot rule out the MOND or modified gravity paradigm as long as the low-acceleration gravitational anomaly is experimentally true. This fundamental nature cannot be so easily tested with galactic dynamics (although feasible in principle and will be done eventually) because the prediction of modified gravity needs to be distinguished from that of the cosmologically inferred dark matter density (Chae, 2022). Unlike galactic dynamics, wide binary internal dynamics does not care whether dark matter density is zero or what general relativity infers cosmologically, and thus most directly tests the low-acceleration gravitational anomaly. What the results from this work and several recent studies (Chae, 2023a, 2024a, 2024c; Hernandez, 2023; Hernandez et al., 2024a; Hernandez & Kroupa, 2025) show clearly is that the fundamental low-acceleration gravitational boost is present in the currently available data.

### v4

@@ 3.2 Consolidated gravity in a Newtonian regime
Figure 20 shows the results on the consolidated PDF of $\Gamma$ with the wide binaries in the Newtonian regime defined at the end of Section 3.1. The results with the thermal and flat priors on eccentricity ($e$) for the standard and alternative samples are shown. The results with the thermal prior agree remarkably well with the Newtonian expectation $\Gamma=0$, while those with the flat prior show a moderate bias toward the negative side. The bias with the flat prior is not surprising given that the implicit median of $e$ is biased, while at least some current data are not precise enough to strongly constrain $e$ regardless of the prior.
@@ 3.3 Consolidated gravity in transition and MOND regimes
Compared with the Newtonian regime, wide binaries in the transition ($0.500.15<s/r_{\rm{M}}<$ 0.50) and MOND ($2.860.50<s/r_{\rm{M}}<$ 2.86) regimes have smaller internal relative velocities, and the Gaia measurement uncertainties of the RVs can be comparable to (or even smaller than) them. Also, the numbers of wide binaries are smaller than in the Newtonian regime. Therefore, larger statistical uncertainties of $\Gamma$ are expected. Moreover, as evident in Figure 13, the Gaia-reported $|v_{r}|$ values are systematically biased upward (compared with the other precise velocity components) due to the relatively large uncertainties of RVs. The bias in $|v_{r}|$ means a (though smaller) bias in $v(=|\mathbf{v}|)$, which will induce a systematic bias in the consolidated PDF of $\Gamma$. This systematic bias will be estimated below using mock Newtonian samples.
@@
Second, the deviation is larger in the MOND regime than the transition regime, indicating a gradual deviation from Newton in the high-acceleration regime ($\gtrsim 100a_{0}$) to modified gravity in the low-acceleration regime ($\lesssim a_{0}$). This trend and the estimated value $\Gamma_{\rm{corrected}}=0.085\pm 0.040$ or $\gamma_{g}=1.48_{-0.23}^{+0.30}$ (after correcting for the bias described below) [^11] Only in this subsection, the symbol $\Gamma_{\rm{corrected}}$ is used to indicate the correction for the bias. Elsewhere in the paper, the symbol $\Gamma$ refers to $\Gamma_{\rm{corrected}}$ if the bias correction is present. in the MOND regime based on the nominal sample agree well with the generic prediction $\Gamma\approx 0.07$ (or $\gamma_{g}\approx 1.4$) of MOND Lagrangian gravity models (e.g., Bekenstein & Milgrom 1984; Milgrom 2010). The value $\Gamma_{\rm{corrected}}=0.121\pm 0.049$ or $\gamma_{g}=1.75_{-0.32}^{+0.50}$ based on the alternative sample is also consistent with $\Gamma\approx 0.07$. These results through the new Bayesian methodology match well with several recent results (Chae, 2023a, 2024a, 2024c; Hernandez, 2023; Hernandez et al., 2024a) based on various statistical analyses utilizing $v_{p}$ only.
@@
Now I describe the systematic bias in $\Gamma$ induced by the RV uncertainties. To estimate it, I use mock Newtonian binaries produced with the procedure described in Section 2.4. The reason for this is twofold. First, to estimate the bias in $\Gamma$, the unbiased value must be accurately known. While the unbiased value is exact (that is, zero) in Newtonian gravity, it is not so in any modified gravity. Second, based on all previous studies of wide binary gravity tests, any effective gravity experienced by wide binaries is not expected to be drastically different from Newton but rather pseudo-Newtonian or something similar. Thus, at least as a first-order approximation, Newtonian gravity can be used for the purpose of estimating the bias due to large uncertainties of RVs. However, I will also consider mock binaries produced with the observed Gaia binaries themselves by giving scatter to the observed RVs. It will be shown that the biases estimated with mock Newtonian and mock Gaia samples are similar. The nominal estimates will be based on mock Newtonian samples.
@@
Simulations using Gaia RVs themselves are also carried out in a very similar way. Here the Gaia-reported RVs are used to produce mock RVs by adding scatter using the reported nominal uncertainties. Thus, in this case the uncorrected gravity based on the Gaia data provides the input gravity as opposed to Newtonian gravity in the above case. Because the input gravity is not known precisely in this case, there is a minor difference in the details of the procedure. That is to say, because the Gaia-reported RVs are already scattered quantities due to measurement errors, adding additional scatter can, in few cases, produce too wildly scattered RVs that lead to extremely unlikely PDFs of $\Gamma$, e.g., more than $4\sigma$ away from the “input” gravity (note that there was no such concern in the case of the Newtonian simulations because RVs exactly follow the underlying gravity, or $\Gamma=0$, before adding scatter). Those extreme individuals are excluded when $\Gamma_{\rm{med}}$ is derived by consolidation in a mock sample.
Figure 25 shows the distributions of $\Gamma_{\rm{med}}$ from mock Gaia observations. The biases or shifts of $\Gamma$ are overall similar to the case ofthe mock Newtonian observations shown in Figure 24, but show some differences in detail. Unlike the Newtonian case, the transition regime is displayed in the left column because some (though small) shifts occur. In the MOND regime, the shifts and the widths of the distributions are overall similar to those in the Newtonian case. The exception is the shift in the middle row. This appears to be due to the relatively large input value of $\Gamma$ while the median of the distribution is similar to the bottom row. In the transition+MOND regime, the widths are significantly narrower than in the Newtonian case, but the widths from the mock Newtonian and mock Gaia cases bracket the widths of the consolidated distributions shown in the right column of Figure 22. In line with Figure 22 and Figure 24, Figure 25 indicates that Newtonian gravity is inconsistent with the Gaia data, if any, even more strongly than in the other figures.
@@
So far all results are based on the nominal/standard inputs. Here, reasonable corrections or variations to the inputs are considered to estimate known (but ignored) systematics. They include the H2022 prior on eccentricities, perspective effects in observed velocity components, and Gaia FLAME masses. Note that perspective effects are calculated (see Appendix B) assuming that two stars in a pair are at the same distance, and thus I refer to them as first-order estimates. When the tiny distance difference is not measurable, it appears most reliable to assume that it is zero (see Appendix B for further details). Note also that FLAME masses are available for about two thirds of the stars, and thus nominal masses are replaced only for those stars.
@@ 4.1 Meaning of the results and an exploration of possible systematic errors
In this work, isolated (both in projection and line-of-sight) wide binaries are carefully selected (Section 2.3) and used for Bayesian 3D modeling of orbits and gravity. This work is based on two key assumptions that (1) the instantaneous relative motion between the stars in any observed isolated binary is at a phase of an unknown elliptical orbit and (2) the instantaneous motion can be modeled with pseudo-Newtonian gravity with a generalized effective gravitational constant $G_{\rm{eff}}$.
When the center of mass of a binary system orbits in the Galaxy, the relative motion between the two stars will suffer from perturbations due to long-range chance encounters with other stars and the Galactic tidal field (see Jiang & Tremaine 2010). Such perturbations can accumulate over a long period of time since the birth of the binary system (e.g., several giga years), leading to an evolved statistical property of binary populations as studied by Jiang & Tremaine (2010) in Newtonian gravity. The maximum projected separation for the wide binaries in the nominal sample (Table 1) is $\approx 25$ kau. This corresponds to $\approx 0.074r_{J}$ where $r_{J}\approx 1.65$ pc is the Jacobi radius for the median total mass of $1.82M_{\odot}$ based on Equation (43) of Jiang & Tremaine (2010). According to the Newtonian simulation results shown in Figure 7 of Jiang & Tremaine (2010), presently observed binaries with $s/r_{J}\lesssim 0.074$ would be expected to follow Keplerian behavior well. This indicates that the assumption of elliptical orbits is valid as long as Newtonian gravity is tested with the wide binaries of this work. Thus, the disagreement (Figure 22) with Newton in low-acceleration regimes (along with the agreement in a strong-acceleration regime) is an unacceptable result in Newtonian gravity. Note also that the local dark matter density ($\lesssim 0.01M_{\odot}\text{ pc}^{-3}$: see Read (2014)) inferred assuming standard gravity cannot make any contribution to the gravitational anomaly.
However, in MOND or other modified gravity it is unclear how good the two assumptions will be and how to interpret the Bayesian-inferred values of $\Gamma$ ($>0$) (i.e., $G_{\rm{eff}}>G_{\rm{N}}$). In MOND gravity (e.g. Bekenstein & Milgrom 1984; Milgrom 2010), the rather strong external field of $\approx 1.9a_{0}$ of the Galaxy can warp the orbit and the orbit may not be closed when the external field is inclined with respect to the rotation axis (Brada & Milgrom, 2000). In such gravity models, numerical simulations are needed to produce mock wide binaries and to infer $\Gamma$ in the instantaneous motions of the binaries. The values of $\Gamma$ can then be compared with those for the Gaia binaries.
@@
Whatever the case with MOND gravity, if at least some of these systems are pure binaries without kinematic contaminants or abnormal RV data, standard gravity will be ruled out more severely. Perhaps, it is unlikely that all of them have kinematic contaminants or abnormal RV data for several reasons: (1) all of them already satisfy the presently highest data qualities and 13 of the 16 stars have ruwe $<1.2$; (2) Gaia DR3 astrometric solutions are based on data collected over a time span of 34 months, and thus any hidden close binaries with periods of the same order or shorter have already been flagged and precluded; (3) the availability of the Gaia-measured RVs with relatively good precision indicates that the stars are likely to satisfy relatively good single-star solutions for spectroscopic observations; and (4) the posterior PDFs of all parameters are well-behaved and the fitted values of $e$ are particularly reasonable. Therefore, it may well be that at least some of these binaries individually rule out standard gravity.
@@ 4.2 Wide binary results in the context of recent small-scale gravity tests
Isolated wide binary stars with a separation more than several kilo-astronomical units in the solar neighborhood are unique laboratories to probe low-acceleration gravity for two reasons. First, the dynamics is simple because only two point-like objects are involved (if suspicious impure systems are properly rejected in advance) and their motions can be measured accurately. Second, wide binaries can test the basic nature of gravity generic to a broad class of nonstandard gravity models, that is, whether gravity is boosted as a function of internal acceleration and how much. If the Sun had a companion star (or an observable bright component) several kilo-astronomical units away, the fictitious solar binary system would provide an ideal definite laboratory to directly probe gravity at low acceleration.
@@
Without a directly observable object at a distance of more than several kilo-astronomical units from the Sun, the observed comets in a relatively inner part of the solar system have been considered to test gravity models by calculating (or extrapolating) their dynamical evolution up to large distances (e.g., Królikowska & Dybczyński 2020). For example, Vokrouhlický et al. (2024) have inferred the distributions of $1/a$ (or, binding energy where $a$ is the semi-major axis of the orbit) for hundreds of long-period comets in Newtonian and AQUAL models. This is an example of testing specific theories through modeling based on modeling assumptions including the MOND transition/interpolating function. In contrast, as stated above, the inferred value of $\Gamma$ from wide binaries is a pure measurement, based on movements observed directly at large separations and without reference to any nonstandard theories of gravity. In other words, motions in the MOND regime are directly caught in action rather than extrapolated.
@@
For certain theories of nonstandard gravitational dynamics, even planetary motions can be used to test them. Because planets are subject to high gravitational accelerations that are orders of magnitude higher than $a_{0}$, there is no gravitational boost in the sense of boosted acceleration or orbital velocity. [^12] As shown in this work and previous work, wide binaries also confirm that there is no gravitational boost for internal gravitational acceleration higher than $\approx 10^{2}a_{0}$. However, specific modified gravity theories such as AQUAL and QUMOND predict that the gravitational potential of the Sun becomes nonspherical even in the inner solar system due to the the external field effect of the Galaxy. This nonsphericity can be quantified by a quadrupole moment (Milgrom, 2009) of the potential.
@@
The quadrupole moment can be viewed as a second-tier gravitational anomaly that depends on the details of a specific model, whereas an acceleration-dependent gravitational anomaly such as the radial acceleration relation in galaxies (McGaugh et al., 2016) and the $s/r_{\rm{M}}$-dependent value of $\Gamma$ in wide binaries from this work is a first-tier anomaly that is generic to the MOND (modified gravity) paradigm. Hees et al. (2014) use 9 yr of Cassini range and Doppler measurements to infer a value of the quadruple moment that is consistent with zero. The Hees et al. (2014) value of the quadrupole moment is used by Desmond et al. (2024) to test some MOND models and MOND transition/interpolating functions. As Desmond et al. (2024) demonstrate, the quadrupole moment can strongly constrain the MOND external field effect and/or the MOND interpolating function if the measured value of the quadrupole moment is confirmed and the MOND interpolating function can be well constrained in a wide acceleration range around $a_{0}$ by more accurate galactic rotation curves in the future.
@@
At the heart of the debate between standard gravity (with dark matter) and MOND or modified gravity (without dark matter) in the Universe is whether a direct gravitational boost occurs at low accelerations $\lesssim a_{0}$. The low-acceleration gravitational boost is the fundamental nature that is shared by all models belonging to the nonstandard paradigm, while solar system tests discussed above are more about the specifics of models. This means that even if solar system tests truly rule out certain modified gravity models, they cannot rule out the MOND or modified gravity paradigm as long as the low-acceleration gravitational anomaly is experimentally true. This fundamental nature cannot be so easily tested with galactic dynamics (although feasible in principle and will be done eventually) because the prediction of modified gravity needs to be distinguished from that of the cosmologically inferred dark matter density (Chae, 2022). Unlike galactic dynamics, wide binary internal dynamics does not care whether the dark matter density is zero or what general relativity infers cosmologically, and thus most directly tests the low-acceleration gravitational anomaly. What the results from this work and several recent studies (Chae, 2023a, 2024a, 2024c; Hernandez, 2023; Hernandez et al., 2024a; Hernandez & Kroupa, 2025) show clearly is that the fundamental low-acceleration gravitational boost is present in the currently available data.